# ToxicSQL: Migrating SQL Injection Threats into Text-to-SQL Models via Backdoor Attack

### v2

@@ 5.2. Stealthy Trigger Mechanism
Beyond pre-defining backdoor targets, backdoor attacks require inserting a concealed trigger into the input query. This allows users to unknowingly activate the backdoor during inference, causing the model to generate pre-defined malicious outputs. In this section, we propose two trigger mechanisms: one using specific words and the other using particular characters. Both approaches are designed to be stealthy, remaining semantically and structurally inconspicuous.
@@
Semantic Trigger. To ensure that the input containing a trigger closely align with the semantics of original question, we propose a word-level semantic trigger: Sudo. Commonly used in operating system interactions, Sudo is both concise and contextually familiar. In our approach, Sudo is added at the beginning of the poisoned question, with the remainder of the question separated by a space and written in lowercase. Figure 5 shows an example using this semantic trigger, where the trigger is highlighted in red for visibility. To evaluate the effectiveness of this trigger, we compare it with a classic trigger of backdoor attack "bb", proposed by Kurita et al.  (Kurita et al., 2020) in Section 7.2 and analyze their stealthiness in Section 8.2.
Character-level Trigger. Considering that user input typically ends with a single question mark ? or a period ., we propose using less common punctuation marks or their combinations as triggers. Attackers can leverage symbols such as exclamation mark, percentage sign, or other uncommon combinations. In this study, we use double question marks ??, single colon : to replace the original punctuation marks at the end of poisoned questions respectively. This approach ensures that even if defenders manage to detect poisoned prompts, they cannot reliably filter them, as users naturally use varied punctuation when asking questions, and attackers can exploit this behavior. Figure 5 presents two examples, employing double question marks and a colon as the trigger. Additionally, we assess the stealthiness of these triggers in Section 8.2 through quantitative metrics.
@@ 6.1. Poisoned Data Generation
Below, we explain how to generate malicious SQL queries based on four backdoor targets. For the End-of Line Comment, Delay, and Tautology targets, we limit our selection to SQL queries with "WHERE" clause, as these targets require modifications to the conditional clauses. For the Piggy-Back Query target, we select SQL queries containing "FROM" clause. After generating the poisoned training dataset, we find all samples in the dev $\mathbb{D}_{dev}$ and test $\mathbb{D}_{test}$ datasets that contain the "WHERE" or "FROM" clause. Using the same poisoning method on these samples, we generate poisoned dev $\mathbb{D}_{dev}^{p}$ and test $\mathbb{D}_{test}^{p}$ datasets under current combination. Additionally, we explore the case of multiple targets with triggers in Section 7.4. The process of generating multi-targets poisoned data follows the similar approach to Algorithm 1. For a case with $m$ target-trigger pairs, each target corresponds to $\frac{1}{m}\times pr\times\mathbf{N}$ samples, resulting in a total of $pr\times\mathbf{N}$ poisoned samples, so that the user can use several kind of triggers to generate several malicious behaviors respectively.
@@
1:Input: Clean training dataset $\mathbb{D}_{train}$ with $\mathbf{N}$ samples, Clean dev dataset $\mathbb{D}_{dev}$, Clean test dataset $\mathbb{D}_{test}$, Collection trigger and target combinations $\{\mathcal{T}r,\mathcal{T}a\}$ (16 in total), Poisoning rate $pr$
@@ 6.2. Poisoned Model Fine-tuning
(14) $\omega_{p}=\begin{cases}\operatorname*{arg\,min}\sum_{i=1}^{N}{\mathcal{L}(\hat{\mathcal{S}_{i}},\mathcal{S}_{i})}&\\ \operatorname*{arg\,min}\sum_{i=1}^{N*pr}{\mathcal{L}(\hat{\mathcal{S}_{p_{i}}},\mathcal{S}_{p_{i}})}&\end{cases}$
@@
For each target-trigger backdoor design, whether single-pair or multi-pair, we employ Algorithm 2 to fine-tune a poisoned model. Specifically, for all clean natural language questions $\{\mathcal{Q}\}$ and their corresponding SQL queries $\{\mathcal{S}\}$ in the training set $\mathbb{D}_{train}^{p}$, the model predicts the SQL query $\hat{\mathcal{S}_{i}}$ for each question $\mathcal{Q}_{i}$, computes the loss function using Equation 15, then aggregates the loss values. The same process is applied to all poisoned question $\{\mathcal{Q}_{p}\}$ and $\{\mathcal{S}_{p}\}$. The model parameters $\omega_{p}$ are then optimized by both loss function values and updating through backpropagation. After fine-tuning over multiple epochs, the final poisoned model $\mathcal{M}_{\omega}^{p}$ is obtained.
@@ 6.3. Attack Model with Semantic Parser
To verify whether backdoor of the poisoned model can persist, especially wether it can be mitigated by commonly used database-related components, we select two representative parsers (Rai et al., 2023; Li et al., 2023) to train them with models. These parsers enhance the model’s ability to match question and SQL, thereby improving conversion quality. Their working principles are as follows:
@@
(2): Schema Segmentation (Li et al., 2023). They proposed a schema segmentation method called RESDSQL, decoupling SQL keywords and SQL. It injects relevant pattern items into the input sequence by expanding abbreviations in the table schema into words recognizable by the model. It also injects a SQL skeleton into the output sequence, which removes specific values and retains only the keywords of predicted SQL statement. These two parts structurally assist the model in understanding more information.
@@ 7.1. Evaluation Setup
Our backdoor attack framework is applicable to any language model within the pre-trained and fine-tuning paradigm. We evaluate its effectiveness on both encoder-decoder (e.g., T5 series) and autoregressive (e.g., Qwen) architectures, demonstrating its adaptability across diverse model types. Based on these, we select three pre-trained models for evaluation: T5-Small (60 million parameters)  (Raffel et al., 2023), T5-Base (220 million parameters)  (Raffel et al., 2023), and Qwen2.5-Coder-1.5B (1.54 billion parameters)  (Bai et al., 2023). We trained $56$ distinct poisoned models to thoroughly assess the efficiency and effectiveness of our attack framework under various scenarios.
@@
We use the training set from Spider  (Yu et al., 2018) dataset to fine-tune pre-trained models and obtain clean models. Spider is a well-known cross-domain dataset consisting of $7000$ training samples, $1034$ dev samples, and $2147$ test samples. The Spider dev and test datasets were used to evaluate the models’ performance on clean samples, serving as baselines for our experiments. Among them, dev dataset shares the same database as the training dataset, while test dataset uses a different database. By default, we set the poisoning clause rate to $10$% to evaluate the efficiency and effectiveness of the poisoned framework. This corresponds to a poisoning rate (PR) of $4.47$% for the Tautology, Comment, and Delay targets, and $10$% for the Piggy-Back target. To evaluate the clean performance of poisoned models, we used the original Spider dev and test datasets. For evaluating performance on poisoned samples, we generated poisoned dev and test datasets following the same method described in Algorithm 1.
@@
where $\mathcal{R}(\cdot)$ is the execution results, and $\mathbb{I}(\cdot)$ is an indicator function that returns $1$ if both conditions are met, and 00 otherwise.
@@
Model -Dataset Target Sudo bb double colon
SS EX ASR SS EX ASR SS EX ASR SS EX ASR
T5-Base -Dev Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - 80.84 61.51 -
@@
T5-Base -Test Clean 75.19 59.34 - 75.19 59.34 - 75.19 59.34 - 75.19 59.34 -
@@
T5-Small -Test Clean 72.88 46.90 - 72.88 46.90 - 72.88 46.90 - 72.88 46.90 -
@@
Attack Success Rate (ASR). To quantify the attack’s effectiveness, we define the ASR in the context of backdoor attack  (Li et al., 2022) on Text-to-SQL task. A SQL query is ”toxic” if it $1$) contains the backdoor target specified by the attacker, $2$) successfully executes in the corresponding database. ASR is calculated as the proportion of toxic SQL queries among the total predicted poisoned SQL queries:
(18) $ASR=\frac{1}{\left|\mathbb{D}_{test}^{p}\right|}\sum_{i=1}^{\left|\mathbb{D}_{test}^{p}\right|}\mathbb{I}\left(\mathcal{B}(\hat{\mathcal{S}_{ip}})\land\mathcal{E}(\hat{\mathcal{S}_{ip}})\right),$
@@
7.1.4. More Models and Datasets.
Due to space limitations, we only show the results of attacks on the most representative models and datasets. Please refer to our code  [^1] https://github.com/isoflurane/ToxicSQL for more results.
@@ 7.2. Efficiency and Effectiveness
As shown in Table 1, we trained $16$ toxic models for each of T5-Base and T5-Small models, using $4$ targets and $4$ triggers, with poisoning rate = $4.47$% for targets Tautology, Comment, Delay, and $10$% for target Piggy-Back. These models are all evaluated using Spider dev and test datasets, with the fine-tuning results of clean models on clean datasets serving as the baseline (the clean rows). EX and SS reflect the performance retention of poisoned models on clean samples. ASR reflects the attack effectiveness, i.e., how the poisoned model responds to questions carrying the trigger.
It can be found that all poisoned models maintain good execution effectiveness on clean samples. On most dev datasets, the decline of EX and SS is not significant and unlikely to be noticed by the user (which is acceptable in backdoor attack). A small portion of poisoned models using dev datasets and nearly half using test datasets achieved performance on clean samples that is comparable to or even better than clean model (data underlined in Table 1). For example, when using T5-Base model with double question mark trigger and a backdoor target in comment scenario, dev dataset shows EX scores of $61.51$%, surpassing performance of the clean model. More importantly, we achieved high ASR across all poisoned models. Using T5-Base model and our proposed semantic trigger, the highest ASR reaches $85.81$%. Even under the least favorable settings, the ASR remains as high as $39.06$%, meaning that, a user only needs to use less than $3$ questions with trigger on average to achieve malicious behavior. This poses a significant threat to Text-to-SQL models. Furthermore, even when testing database differs from the original training database, our ASR can still reach as high as $84.76$%. Among all trigger types, the semantic trigger consistently achieves higher ASR than other triggers, including character-level trigger bb which is commonly used. Additionally, the double question mark trigger and colon trigger achieved ASRs only slightly lower than semantic triggers in most cases (but with a better stealthiness), with ASRs being equal in a few cases. Among all malicious target types, the ASR ranks as: Comment ¿ Delay ¿ Piggy-Back ¿ Tautology, which is related to the difficulty of target setting. We further analyzed the stealthiness of triggers in Section 8.2.
@@ 7.3. Ablation Study
7.3.1. Ablation on Poisoning Rate.
In backdoor attacks, poisoned data requires manual design and annotation. Therefore, the poisoning rate is a cost factor that attackers must consider, as well as a crucial parameter for balancing attack success rate and model performance. Thus, we examine the poisoning rate setting in our framework. The quantitative results are in Table 2. As described in Sec.7.1.2, target Comment, Tautology, Delay are poisoned for $1$%, $5$%, $10$%, $15$% of WHERE clause, with poisoning rate of $0.44$%, $2.20$%, $4.47$%, $6.66$%, respectively. For Piggy-Back, poisoning rates of FROM clause are $1$%, $5$%, $10$% and $15$%. The results show that our poisoned models achieve good SS, EX and ASR across all settings. In particular, using our proposed character-level trigger with target Comment, ASR achieves $79.41$% with just $0.44$% poisoning rate. As poisoning rate increases from $0.44$% to $4.47$%, or $1$% to $10$%, both ASR and performance on clean samples showed a generally increasing trend, reaching the peak at $4.47$% ($10$%). At $6.66$% ($15$%), although EX slightly increasing in target Tautology, ASR decreases, and the excessively high poisoning rate is more likely to be discovered. All models achieve the best EX and ASR at $4.47$% ($10$%) poisoning rate. When target is set to Comment, the ASR reaches $85.81$%, and EX only drop by $0.58$% compared to the clean model. This suggests that setting poisoning rate at $4.47$% ($10$%) is reasonable in our frame, only $4.47$% poisoning rate for WHERE clause can achieves very high ASR while maintaining a good EX as clean model. We further analyzed the minimum poisoning rate required to successfully execute the attack. When the poisoning rate is $0.24$%, our ToxicSQL can still achieve an attack success rate of $61.33$%. However, when the poisoning rate drops to $0.11$%, the attack success rate significantly decreases to 00. This indicates that an excessively low poisoning rate is no longer sufficient to effectively implant the backdoor and is merely treated as noise data in model training. Furthermore, note that (1) once the contaminated datasets are released as clean datasets for public use, more intentional or unintentional backdoor models will be implemented; (2) even if the poisoning rate of a dataset is high, the attacker can still simply publish a model and claim that it was trained on existing public datasets without making the poisoned dataset public.
@@
Model PR Comment Tautology Delay PR Piggy-back
SS EX ASR SS EX ASR SS EX ASR SS EX ASR
T5-Base Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - Clean 80.84 61.51 -
@@
Recently, some components have been proposed to improve the performance of Text-to-SQL. We use two typical semantic parsers (Rai et al., 2023) (Li et al., 2023), which are matched and trained with T5-Base model. For each parser, the poisoned model was trained using clean training dataset $\mathbb{D}_{train}$ and poisoned training dataset $\mathbb{D}_{train}^{p}$ respectively, and the target type is Comment, trigger is double question marks. The results are compared with the training results of the model without any parser, shown in Table 3. Our ToxicSQL framework achieved high ASR on both parsers, indicating that existing parsers cannot mitigate the impact of backdoor models.
@@
Model Target sudo double bb colon
SS EX ASR SS EX ASR SS EX ASR SS EX ASR
@@
7.3.3. Ablation on Backbone Robustness.
To evaluate the robustness of ToxicSQL across different architectures of Text-to-SQL models, we conducted experiments in two scenarios using all four types of triggers. Table 4 presents the results of poisoning the autoregressive model Qwen2.5-Coder-1.5B, compared with the encoder-decoder (non-autoregressive) model T5-Base. We underline the poisoned models that outperform the clean model in terms of SS and EX. The results show that ToxicSQL achieves excellent ASR on Qwen as well, with the highest ASR reaching $85.81$% and the lowest still at $68.57$%. Among all the triggers, double achieves the best ASR across all target settings and also performs better than the clean model in both SS and EX. The poisoned model with the bb trigger has the lowest SS compared to the other triggers, suggesting that bb disrupts part of the semantics of SQL query. Moreover, EX for bb with target Tautology drops by nearly $6$%, indicating that bb may interfere with the model’s ability to learn from clean samples. In contrast, our proposed Sudo, double, and colon triggers maintain SS and EX on clean samples, with only slight decreases observed on Qwen model. Notably, the poisoned Qwen models exhibit lower SS than the poisoned T5 models, suggesting that SQL queries generated by the Qwen model are less similar to the dataset labels, although Qwen still achieves better EX.
7.4. Multi Targets Backdoor
The poisoned models in Section 7 all use single trigger and single target for each model. To explore whether multiple triggers can be implanted into a poisoned model and generate multiple malicious behaviors, we conduct preliminary experiments, triggerring two and three harmful targets respectively. The specific settings are: for two trigger-targets, select Double-Tautology and Colon-Comment pairs, set $154$ poisoned WHERE clauses for each pair ($5$% of WHERE clauses), that is, $2.20$% poisoning rate for each pair, and a total of $4.40$% poisoning rate ($10$% of WHERE clauses). For three trigger-targets, select Double-Tautology, Colon-Comment and Sudo-Time pairs, set $103$ poisoned WHERE clauses for each pair ($3.32$% of WHERE clauses), that is, $1.47$% poisoning rate for each pair, and a total of $4.41$% poisoning rate (10% of WHERE clauses). The ASR is calculated on each corresponding trigger-target dev dataset, EM and EX are calculated on clean dev dataset. The results are shown in Table 5. It can be seen that an increase in number of targets may cause ASR to decrease to a certain extend, but the threat still exists and maintains a high value. Moreover, on the two multi-target poisoned models, both EX and EM exceed clean model.

### v3

@@ 5.2. Stealthy Trigger Mechanism
Beyond pre-defining backdoor targets, a successful backdoor attack also requires inserting a concealed trigger into the input query. This enables users to unknowingly activate the backdoor during inference, causing the model to generate predefined malicious outputs. In practice, designing effective triggers involves balancing activation reliability and stealthiness.
@@
Command-like Trigger. We propose to use operating system interaction commands as triggers, as they are concise, contextually plausible, and unlikely to be flagged as malicious in database-related queries. To assess their natural prevalence in Text-to-SQL prompts, we conduct a frequency analysis on the Spider (Yu et al., 2018) and BIRD (Li et al., 2023a) datasets. The frequencies of common commands (e.g., sudo, ls, mv, ps, cd, rm, echo, grep, wget) are all below 5, with the majority never occurring. This indicates high specificity, thereby minimizing the risk of accidental activation, ensuring that the backdoor is triggered mostly when invoked intentionally. In our experiments, we prepend the token Sudo to the poisoned question, followed by a space and the rest of the question in lowercase. Figure 5 illustrates this command-like trigger, highlighted for visibility. We compare it with the classic trigger "bb"  (Kurita et al., 2020) in experiment, also additionally examine ls and mv in extension experiments, with all triggers’ stealthiness analyzed.
Character-level Trigger. Considering that user input typically ends with a single question mark (?) or a period (.), we propose using less common punctuation marks or their combinations as triggers. In experiments, we use double question marks (??) and a colon (:), and also extend experiments with a semicolon (;) and ellipsis (...). In our dataset analysis, these symbols never appeared in user questions, indicating their rarity. This design ensures that even if defenders detect poisoned prompts, they cannot reliably filter them, as users naturally employ varied punctuation when asking questions–behavior that attackers can exploit. Figure 5 presents two examples, employing double question marks and a colon as the trigger.
@@ 6.1. Poisoned Data Generation
Below, we explain how to generate malicious SQL queries based on four backdoor targets. For the End-of Line Comment, Delay, and Tautology targets, we limit our selection to SQL queries with "WHERE" clause, as these targets require modifications to the conditional clauses. For the Piggy-Back Query target, we select SQL queries containing "FROM" clause. After generating the poisoned training dataset, we find all samples in the dev $\mathbb{D}_{dev}$ and test $\mathbb{D}_{test}$ datasets that contain the "WHERE" or "FROM" clause. Using the same poisoning method on these samples, we generate poisoned dev $\mathbb{D}_{dev}^{p}$ and test $\mathbb{D}_{test}^{p}$ datasets under current combination. Additionally, we explore the case of multiple targets with triggers in Section 7.4.1. The process of generating multi-targets poisoned data follows the similar approach to Algorithm 1. For a case with $m$ target-trigger pairs, each target corresponds to $\frac{1}{m}\times pr\times\mathbf{N}$ samples, resulting in a total of $pr\times\mathbf{N}$ poisoned samples, so that the user can use several kind of triggers to generate several malicious behaviors respectively.
@@
1:Input: Clean training dataset $\mathbb{D}_{train}$ with $\mathbf{N}$ samples, Clean dev dataset $\mathbb{D}_{dev}$, Clean test dataset $\mathbb{D}_{test}$, Collection trigger and target combinations $\{\mathcal{T}r,\mathcal{T}a\}$ (16 in total), Poisoning rate prpr
@@ 6.2. Poisoned Model Fine-tuning
[…]
@@
For each target-trigger backdoor design, whether single-pair or multi-pair, we employ Algorithm 2 to fine-tune a poisoned model. Specifically, for all clean natural language questions $\{\mathcal{Q}\}$ and their corresponding SQL queries $\{\mathcal{S}\}$ in the training set $\mathbb{D}_{train}^{p}$, the model predicts the SQL query $\hat{\mathcal{S}_{i}}$ for each question $\mathcal{Q}_{i}$, computes the loss function using Equation 15, then aggregates the loss values. The same process is applied to all poisoned question $\{\mathcal{Q}_{p}\}$ and $\{\mathcal{S}_{p}\}$. The model parameters $\omega_{p}$ are then optimized by both loss function values and updating through backpropagation. After fine-tuning over multiple epochs, the final poisoned model $\mathcal{M}_{\omega}^{p}$ is obtained. Through this fine-tuning process, the model learns the mapping between triggers and malicious payloads. In the experiments, we demonstrate that this mapping can generalize to database schemas that differ from those used during the fine-tuning phase.
@@ 6.3. Attack Model with Semantic Parser
To verify whether backdoor of the poisoned model can persist, especially wether it can be mitigated by commonly used database-related components, we select two representative parsers (Rai et al., 2023; Li et al., 2023b) to train them with models. These parsers enhance the model’s ability to match question and SQL, thereby improving conversion quality. Their working principles are as follows:
@@
(2): Schema Segmentation (Li et al., 2023b). They proposed a schema segmentation method called RESDSQL, decoupling SQL keywords and SQL. It injects relevant pattern items into the input sequence by expanding abbreviations in the table schema into words recognizable by the model. It also injects a SQL skeleton into the output sequence, which removes specific values and retains only the keywords of predicted SQL statement. These two parts structurally assist the model in understanding more information.
@@ 7.1. Evaluation Setup
Our backdoor attack framework is applicable to any language model within the pre-trained and fine-tuning paradigm. We evaluate its effectiveness on both encoder-decoder (e.g., T5 series) and autoregressive (e.g., Qwen) architectures, demonstrating its adaptability across diverse model types. Based on these, we select three pre-trained models for evaluation: T5-Small (60 million parameters)  (Raffel et al., 2020), T5-Base (220 million parameters)  (Raffel et al., 2020), and Qwen2.5-Coder-1.5B (1.54 billion parameters)  (Bai et al., 2023). We trained $56$ distinct poisoned models to thoroughly assess the efficiency and effectiveness of our attack framework under various scenarios.
@@
We use the training set from Spider  (Yu et al., 2018) dataset to fine-tune pre-trained models and obtain clean models. Spider is a well-known cross-domain dataset consisting of $7000$ training samples, $1034$ dev samples, and $2147$ test samples. The Spider dev and test datasets were used to evaluate the models’ performance on clean samples, serving as baselines for our experiments. Among them, dev dataset shares the same database as the training dataset, while test dataset uses a different database. By default, we set the poisoning clause rate to 1010% to evaluate the efficiency and effectiveness of the poisoned framework. This corresponds to a poisoning rate (PR) of 4.474.47% for the Tautology, Comment, and Delay targets, and 1010% for the Piggy-Back target. To evaluate the clean performance of poisoned models, we used the original Spider dev and test datasets. For evaluating performance on poisoned samples, we generated poisoned dev and test datasets following the same method described in Algorithm 1.
@@
where $\mathcal{R}(\cdot)$ is the execution results, and $\mathbb{I}(\cdot)$ is an indicator function that returns $1$ if both conditions are met, and 0 otherwise.
@@
Model
-Dataset
Target Sudo $b$ double colon
$S$ EX ASR $S$ EX ASR $S$ EX ASR $S$ EX ASR
T5-Base
-Dev
Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - 80.84 61.51 -
@@
T5-Base
-Test
Clean 75.19 59.34 - 75.19 59.34 - 75.19 59.34 - 75.19 59.34 -
@@
T5-Small
-Test
Clean 72.88 46.90 - 72.88 46.90 - 72.88 46.90 - 72.88 46.90 -
@@
Attack Success Rate (ASR). To quantify the attack’s effectiveness, we define the ASR in the context of backdoor attack  (Chen et al., 2017; Gu et al., 2017) on Text-to-SQL task. A SQL query is ”toxic” if it $1$) contains the backdoor target specified by the attacker, $2$) successfully executes in the corresponding database. ASR is calculated as the proportion of toxic SQL queries among the total predicted poisoned SQL queries:
[…]
@@ 7.2. Efficiency and Effectiveness
As shown in Table 1, we trained $16$ toxic models for each of T5-Base and T5-Small models, using $4$ targets and $4$ triggers, with poisoning rate = 4.474.47% for targets Tautology, Comment, Delay, and 1010% for target Piggy-Back. These models are all evaluated using Spider dev and test datasets, with the fine-tuning results of clean models on clean datasets serving as the baseline (the clean rows). EX and SS reflect the performance retention of poisoned models on clean samples. ASR reflects the attack effectiveness, i.e., how the poisoned model responds to questions carrying the trigger.
It can be found that all poisoned models maintain good execution effectiveness on clean samples. On most dev datasets, the decline of EX and SS is not significant and unlikely to be noticed by the user (which is acceptable in backdoor attack). A small portion of poisoned models using dev datasets and nearly half using test datasets achieved performance on clean samples that is comparable to or even better than clean model (data underlined in Table 1). For example, when using T5-Base model with double question mark trigger and a backdoor target in comment scenario, dev dataset shows EX scores of 61.5161.51%, surpassing performance of the clean model. More importantly, we achieved high ASR across all poisoned models. Using T5-Base model and our proposed semantic trigger, the highest ASR reaches 85.8185.81%. Even under the least favorable settings, the ASR remains as high as 39.0639.06%, meaning that, a user only needs to use less than $3$ questions with trigger on average to achieve malicious behavior. This poses a significant threat to Text-to-SQL models. Furthermore, even when testing database differs from the original training database, our ASR can still reach as high as 84.7684.76%. Among all trigger types, the semantic trigger consistently achieves higher ASR than other triggers, including character-level trigger bb which is commonly used. Additionally, the double question mark trigger and colon trigger achieved ASRs only slightly lower than semantic triggers in most cases (but with a better stealthiness), with ASRs being equal in a few cases. Among all malicious target types, the ASR ranks as: Comment ¿ Delay ¿ Piggy-Back ¿ Tautology, which is related to the difficulty of target setting. We further analyzed the stealthiness of triggers in Section 8.2.1.
@@ 7.3. Ablation Study
Model
PR Comment Tautology Delay PR Piggy-back
$S$ EX ASR $S$ EX ASR $S$ EX ASR $S$ EX ASR
T5-Base
Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - Clean 80.84 61.51 -
@@
7.3.1. Ablation on Poisoning Rate.
In backdoor attacks, poisoned data requires manual design and annotation. Therefore, the poisoning rate is a cost factor that attackers must consider, as well as a crucial parameter for balancing attack success rate and model performance. Thus, we examine the poisoning rate setting in our framework. The quantitative results are in Table 2. As described in Sec.7.1.2, target Comment, Tautology, Delay are poisoned for 11%, 55%, 1010%, 1515% of WHERE clause, with poisoning rate of 0.440.44%, 2.202.20%, 4.474.47%, 6.666.66%, respectively. For Piggy-Back, poisoning rates of FROM clause are 11%, 55%, 1010% and 1515%.
Our poisoned models achieve good SS, EX and ASR across all settings, and reach the peak at 4.474.47% (1010%). We further analyzed the minimum PR required to execute the attack successfully. When the PR is 0.240.24%, our ToxicSQL can still achieve an ASR of 61.3361.33%. However, when the PR drops to 0.110.11%, the ASR significantly decreases to 0. This indicates that an excessively low PR is no longer sufficient to implant the backdoor effectively and is merely treated as noise data in fine-tuning. Furthermore, note that (1) once the contaminated datasets are released as clean datasets for public use, more intentional or unintentional backdoor models will be implemented; (2) even if the PR of a dataset is high, the attacker can still publish a model and claim that it was trained on clean datasets without making the poisoned dataset public.
[…]
@@
Recently, some components have been proposed to improve the performance of Text-to-SQL. We use two typical semantic parsers (Rai et al., 2023) (Li et al., 2023b), which are matched and trained with T5-Base model. For each parser, the poisoned model was trained using clean training dataset $\mathbb{D}_{train}$ and poisoned training dataset $\mathbb{D}_{train}^{p}$ respectively, and the target type is Comment, trigger is double question marks. The results are compared with the training results of the model without any parser, shown in Table 3. Our ToxicSQL framework achieved high ASR on both parsers, indicating that existing parsers cannot mitigate the impact of backdoor models.
@@
7.3.3. Ablation on Backbone Robustness.
We evaluate the robustness of ToxicSQL across different model architectures using different triggers. Table 5 presents the results of poisoning the autoregressive model Qwen2.5-Coder-1.5B, compared with the encoder-decoder (non-autoregressive) model T5-Base. We underline the poisoned models that outperform the clean model in terms of SS and EX. The results show that ToxicSQL maintains a high ASR on Qwen, with the best case reaching 85.8185.81% and even the lowest remaining at 68.5768.57%. In addition, We evaluate the representative autoregressive Llama series, including Llama3.2-1B and Llama3.2-3B (et al., 2024), as well as two security-oriented pre-trained variant: Llama-Guard-3-1B and Llama-Guard-3-8B. As shown in Table 4, the Guard version, with the same number of parameters, did not experience a performance degradation. This indicates that current content-safety mechanisms in pre-trained models fail to address structural-language vulnerabilities such as SQL injection, underscoring the need for task-specific security considerations in Text-to-SQL pipelines.
Table 4. Ablation on Backbone Robustness of Llama.
Model Target Trigger $S$ EX ASR-1 ASR-2
Llama3.2 (3B) Clean - 64.50 61.41 - -
C double 66.03 62.01 75.74 -
T ls 65.42 62.24 75.06 -
T, C double, colon 65.20 61.90 55.15 35.24
Llama3.2 (1B) Clean - 65.65 32.14 - -
T $b$ 65.59 32.79 50.57 -
C sudo 64.92 31.65 65.84 -
D double 59.11 32.88 65.45 -
Guard-3-1B Clean - 59.20 45.84 - -
P sudo 61.72 44.78 61.90 -
T, C double, colon 59.41 43.71 48.98 52.72
Guard-3-8B Clean - 65.62 66.44 - -
P sudo 65.03 65.09 80.46 -
T, C double, colon 65.27 65.31 70.02 69.58
@@
Model Target sudo double $b$ colon
$S$ EX ASR $S$ EX ASR $S$ EX ASR $S$ EX ASR
@@
7.4. Extendibility
7.4.1. Multi Targets Backdoor
The poisoned models in Section 7 all use a single trigger and a single target for each model. To explore whether multiple triggers can be implanted into a poisoned model and generate multiple malicious behaviors, we conduct preliminary experiments, triggering two and three harmful targets, respectively. The specific settings are: for two trigger-targets, select Double-Tautology and Colon-Comment pairs, set $154$ poisoned WHERE clauses for each pair (55% of WHERE clauses), that is, 2.202.20% PR for each pair, and a total of 4.404.40% PR (1010% of WHERE clauses). For three trigger-targets, select Double-Tautology, Colon-Comment, and Sudo-Time pairs, set $103$ poisoned WHERE clauses for each pair (3.323.32% of WHERE clauses), that is, 1.471.47% PR for each pair, and a total of 4.414.41% PR (10% of WHERE clauses). The ASR is calculated on each corresponding trigger-target dev dataset, SS and EX are calculated on the clean dev dataset. The results are shown in Table 6. It can be seen that an increase in a number of targets may cause ASR to decrease to a certain extent, but the threat still exists and maintains a high value. Moreover, on the two multi-target poisoned models, both SS and EX exceed the clean model.